OpenStudy (anonymous):

1/4x^2-3x+3=0 solve using the quadratic formula

OpenStudy (anonymous):

quadratic formula is -b+or- the square root of b^2-4ac all over 2a. a in this case is 1/4, b is -3, c is 3.

OpenStudy (anonymous):

\[-(-3)\pm \sqrt{(-3)^{2}-4(1/4)(3)}\div2(1/4)\]

OpenStudy (anonymous):

why is b -3

OpenStudy (anonymous):

which is positive 3 since the negative of a negative is a positive plus or minus the square root of negative 3 squared, which is positive 9, since a negative times a negaitve is a positive minus 4 times 1/4, which is 1, times 3 is 3. so inside the radical you have

OpenStudy (anonymous):

cant you multiply everything by 4 to get rid of the fraction

OpenStudy (anonymous):

multiply by 4 to give yourself a break with all these fractions

OpenStudy (anonymous):

9-3 which is 6 and that is now 3 plus or minus the sqare root of 6 all over 2 times 1/4 which is 1/2

OpenStudy (anonymous):

ok and when i multiply by 4 i get x^2-12x+12=0

OpenStudy (anonymous):

\[x^2-12x+12=0\] then easy to use formula. easier still to complete the square: \[(x-6)^2=-12+36=24\] \[x-6=\pm\sqrt{24}\] \[x=6\pm\sqrt{24}\]

OpenStudy (anonymous):

Yes, it's so much easier to complete the square. That's what satellite taught me in precalc, and now my classmates all use quad formula, and I solve the answer much faster.

OpenStudy (anonymous):

no muss no fuss. and you can even write s \[x=6\pm 2\sqrt{6}\] if you like

OpenStudy (anonymous):

ok i got another problem x^2+7x+2=0

OpenStudy (anonymous):

if i complete the square i get (x+49/4)^2